OpenStudy (lukecrayonz):

How to differentiate? x^2/x-11 Quotient rule?

ganeshie8 (ganeshie8):

u dont like quotient rule, do u

zepdrix (zepdrix):

lol :)

ganeshie8 (ganeshie8):

he likes converting it to neg exponents and working product rule i think :)

OpenStudy (lukecrayonz):

^ Indeed

zepdrix (zepdrix):

Product Rule is easier to remember, that's for sure. Don't have to worry about getting the subtraction backwards. Quotient Rule can save you some trouble with simplifying though :o

OpenStudy (lukecrayonz):

Okay quickly answer this question and then get back to the first question: f(u)=u^2 g(x)=u=x^5+2 Find (f o g)'(-1) Would it just be f'(-1)=(x^5+2)^2

zepdrix (zepdrix):

\[\Large \left(\frac{x^2}{x-11}\right)'\quad=\quad \frac{\color{royalblue}{(x^2)'}(x-11)-x^2\color{royalblue}{(x-11)'}}{(x-11)^2}\]

zepdrix (zepdrix):

That's for the first question. Oh composition stuff? :hmm

zepdrix (zepdrix):

g(x)=u ??? Hmm so I guess we're taking our derivative with respect to uhhhhhhh...... x i think.\[\Large f(\color{teal}{g(x)})\quad=\quad f(\color{teal}{u}) \quad=\quad u^2 \quad=\quad (x^5+2)^2\]Mmmm I dunno, this one is so weird... D:

zepdrix (zepdrix):

@hartnn

zepdrix (zepdrix):

\[\Large \frac{d}{dx}f(g(x))\quad=\quad 2(x^5+2)(5x^4)\] Mmmmmm like that maybe? +_+

zepdrix (zepdrix):

Evaluated at x=-1,\[\Large 2\left[(-1)^5+2\right]\left[5(-1)^4\right]\]

zepdrix (zepdrix):

Somethinggggggg like that? :o